PLEASE HELP ME SOLVE THIS PROBLEM 12<= 3/2 X + 6
Answer by checkley71(8403)

12<=3/2x+6 or 12-6=3/2x or 6=3/2x or 12=3x or x=12/3 or x=4
prppf 12=3/2*4+6 or 12=12/2+6 or 12=12